OpenStudy (kammykoo):

The axis of symmetry for the graph of the function is f(x) = x2 + bx + 10 is x = 6. What is the value of b?

OpenStudy (lapinblanc):

-12 The axis of symmetry is a vertical line x=(-b)/(2a) You know the following: x=6 a=1 (because x squared has no coefficient, therefore is automatically 1) b=unknown Now you input. 6=(-b)/(2*1) multiply bottom 6=(-b)/2 multiply by reciprocal of denominator 2*6=(-b/2)*(2/1) cross out and multiply 12=-b multiply both sides by -1 to cancel out the -b -12=b so b =-12